OpenStudy (anonymous):

Misha deposited $3,725 into a savings account 22 years ago. The account has an interest rate of 2.2% and the balance is currently $6,012.37. How often does the interest compound?

OpenStudy (anonymous):

monthly quarterly annually daily

OpenStudy (dumbcow):

6012.37 = 3725(1+.022/n)^22n plug in n=1,4,12,365 and see which one works

OpenStudy (anonymous):

i did they all came out with balances in the $8000's :(

OpenStudy (dumbcow):

somethings wrong with your calculation then, they should all be in the 6000's

OpenStudy (dumbcow):

try n =1 again 6012.37 = 3725(1+.022)^22

OpenStudy (anonymous):

oh dear missed that one haha thank you :) your a life saver >.<!